There are 1000 identical individual consumers for a good with a demand function Qd=12-20p and 100 identical suppliers with the supply function Qs=20p, find the demand and supply schedule and also find the equilibrium graphically

1
Expert's answer
2022-01-10T14:43:19-0500

Solution:

Derive market demand and market supply:

Market demand: 1000(12 – 20p) = 12,000 – 20,000P

Market supply: 100(20P) = 2,000P

 

At equilibrium: Qd = Qs

12,000 – 20,000P = 2,000P

12,000 = 2,000P + 20,000P

12,000 = 22,000P

P = 0.55

Equilibrium price = 0.55

Derive quantity:

Qd = 12,000 – 20,000P = 12,000 – 20,000(0.55) = 12,000 – 11,000 = 1,000

Equilibrium quantity = 1,000

 

The graph is as below showing the equilibrium quantity and price: